OpenStudy (anonymous):

Which of the following equations is an example of the general form of the equation of a line in standard form? A. y = mx + b B. Ax = By + C C. By = Ax + C D. Ax + By = C

OpenStudy (anonymous):

its A

OpenStudy (***[isuru]***):

yes! .... wt's ur question then ? u already know the answer.....XD

OpenStudy (anonymous):

lol aye can you help me with a question?

OpenStudy (anonymous):

Write the equation of the line that passes through (–2, 1) and is perpendicular to the line 3x – 2y = 5

OpenStudy (anonymous):

First turn 3x-2y = 5 into standard form (y = mx +b). Then, since you know that a line perpendicular to another line has the negative reciprocal slope of the original line, you know the 2nd line's slope (m). Then you plug in x and y into that 2nd line equation and solve for b (the y-intercept). Once you have be, you go back to your 2nd line equation and plug in b, and bam, you have your equation of the perpendicular line. :) Try this!

OpenStudy (anonymous):

be = b

OpenStudy (anonymous):

(was a typo)

OpenStudy (anonymous):

y=3/2x+4 right?

OpenStudy (***[isuru]***):

when 2 lines become perpendicular to each other , if the slope of one line is "m" , then the slope of the other is "(-1/m)" lets take 3x – 2y = 5 as the first one y = (3/2)x - 5/2 the slope of this one is 3/2 then the slope of the other one will be -2/3 so the equation of the other line will be y = mx + b = (-2/3)x + b now let's consider the coordinates pair (-2 , 1) and substitute those value in our eqation 1 = (-2/3)-2 + b b = 1- 4/3 = -1/3 therefor the equation will be y= (-2/3)x - 1/3 hope this will help ya!